# Automotive Speed Estimation: Sensor Types and Error Characteristics from OBD-II to ADAS

### v2

@@ II Related Work
In addition to built-in vehicular speed sensing technologies and DMIs, alternative systems have been explored for specialized ground vehicle testing. Magnetic Rotary Encoders (MREs) [7] and Non-contact Optical Sensors [8, 9] are examples of technologies used in extreme conditions, such as snow, ice, or wet surfaces, or in harsh environments characterized by dust, moisture, extreme temperatures, and mechanical shocks. Despite their robustness and precision, these niche applications and high costs limit their practical use in everyday automotive systems.
@@
Aspect OBD-II Standard CAN Bus
Purpose Diagnostics and emissions monitoring Real-time communication between ECUs
Standardization Standardized PIDs (e.g., speed: 0x0D) Manufacturer-specific (e.g., CAN ID varies)
Data Access Limited to a subset of vehicle parameters Full access to all CAN messages
Communication Model Request-Response Broadcast (continuous streaming)
Speed of Access Slower, query-based Faster, real-time
Physical Access OBD-II port OBD-II port (or direct access to CAN wires)
In addition to the often omitted derivation process, the navigation and instrumentation literature offers limited clarity on the specific sensors automotive suppliers use for speed estimation. Researchers often overlook the type of speed sensors employed in their test vehicles, commonly employ the general term odometer which covers a wide variety of instruments, and frequently rely on traditional Inertial Measurement Unit (IMU) and GNSS error modeling techniques for sensor fusion [4, 12, 13, 14]. While these approaches have been widely used, they may not guarantee optimal estimation of vehicle speed. Therefore, identifying and characterizing vehicular speed sensors is critical for designing robust and effective navigation systems. Fig. 2 categorizes vehicle speed-sensing technologies into two primary groups: built-in diagnostic data and alternative sensors and systems. The first group includes OBD-II-derived speed measurements and CAN bus-derived data from Vehicle Speed Sensors (VSS) and WSS, both of which are widely integrated into vehicles’ onboard systems. The second group consists of external sensors, such as non-contact optical sensors, DMIs, and perception-based systems like cameras, LiDARs, and RADARs. Cameras and RADARs, in particular, are increasingly utilized not only for ADAS but also for forward velocity estimation. Recent studies have demonstrated the effectiveness of cameras for ego-motion tracking [15] and RADARs for ego-velocity estimation [16]. These perception-based approaches offer practical and cost-effective alternatives, particularly in scenarios where traditional automotive speed sensors are limited or unavailable.
@@ III OBD-II Vehicle Speed Derivation
Modeling errors in consumer-level OBD-II-derived speed measurements presents a significant challenge due to the restricted data available through the vehicle’s regular OBD-II protocol. Generally, the consumer-level OBD-II provides data on the overall vehicle speed rather than the speed of each individual wheel. Additionally, the speed data is often in integer format, which lacks the resolution of decimal points found in double or float types commonly available in CAN bus systems, leading to less precise measurements. When querying the OBD-II system for vehicle speed, a command is sent, and the system responds with encoded data. Below is an example illustrating this process:
@@
The command 010D represents a request for the current vehicle speed (PID 0x0D). The response 41 0D 3C indicates the vehicle speed as 0x3C, which is 60 $km/h$ when converted to decimal [^1] Automakers may configure ECUs to report diagnostic data in region-specific units. For example, vehicle speed may be in miles per hour ($mph$) for the USA or in kilometers per hour ($km/h$) for Canada and Europe.. In other words, the reported speed through OBD-II interfacing is:
@@ III-A Transmission-based Vehicle Speed Determination
The estimation of vehicle speed ($V_{\text{v}}$) from transmission data relies on the relationship between engine speed ($N_{e}$) in RPM and drivetrain parameters, especially gear ratio. The gear ratio governs the conversion of engine speed (input) to wheel speed (output). For example, in Fig. 3, a driving gear with 8 teeth and a driven gear with 16 teeth creates a 2:1 ratio, where the output rotates at half the input speed. This adjustment is crucial for controlling speed and torque. The gearbox alters engine speed based on the engaged gear ratio ($i_{x}$), which changes with each gear shift, while the differential uses its gear ratio ($i_{0}$) to balance wheel speeds, critical for smooth turning.
@@
$V_{\text{v}}=\left\lfloor\left({\frac{N_{e}\cdot\pi\cdot r_{w}}{30\cdot i_{x}\cdot i_{0}}}\right)\right\rceil,$ (2)
$N_{e}$ represents the engine speed, typically measured by a crankshaft position (CKP) sensor or an engine speed sensor. The wheel radius ($r_{w}$), measured in meters ($m$), converts the rotational speed of the wheels into linear speed. This calculation assumes a slip-free mechanical connection between the engine and wheels, ensuring efficient power transfer and accurate speed estimation.
@@ III-B ABS-based Vehicle Speed Determination
where $V_{\text{w}}$ represents the speed of a single wheel in $m/s$, $r_{\text{w}}$ denotes the radius of the wheel in $m$, $N_{p}$ is the count of pulses detected within the time interval $\Delta t$, and $N_{t}$ is the total number of teeth on the reluctor ring. To determine the vehicle’s overall speed, the ECU averages the rotational speeds of all the individual wheels. This average is then used for various control and monitoring functions within the vehicle’s ABS and electronic stability control systems (i.e., ESC).
Compared to the transmission-based system, ABS-based speed estimation offers greater robustness and accuracy while involving fewer parameters. Thus, mathematically, the actual ABS wheel speed reported through the regular consumer-level OBD-II protocol is computed as follows:
$V_{\text{v}}=\left\lfloor\left(\frac{V_{\text{w}_{1}}+V_{\text{w}_{2}}+V_{\text{w}_{3}}+V_{\text{w}_{4}}}{4}\right)\right\rceil,$ (4)
where $V_{\text{v}}$ is the speed of the vehicle, $V_{\text{w}_{i}}$ is the velocity of the $i^{th}$ wheel in $m/s$, assuming each wheel has an installed WSS. To overcome these challenges without relying on automotive OEM collaboration or specialized tools (i.e., reverse engineering the CAN bus) to access detailed information about wheel sensors in test vehicles, navigation, and mapping entities frequently employ DMIs.
@@ III-C Wheel Speed Sensor Types
When dealing with WSS, limited information can be found in the navigation and instrumentation literature related to the type of WSS automotive suppliers use and their corresponding accuracy. Within the automotive industry, two primary types of wheel speed sensors are currently utilized, each possessing unique advantages and disadvantages. These are referred to as $passive$ and $active$ wheel speed sensors, respectively. The operational mechanisms of these sensors exceed the scope of this paper, however, their main differences are shown in Table II.
@@ IV-A Experimental Setup
The experiments were conducted using an ABS-equipped Toyota Sienna minivan, with the logging setup briefly depicted in Fig. 6. The dataset collected during these experiments has been made publicly available [^2] The dataset can be accessed at: https://github.com/hanymragab/navinst19.
@@ IV-B1 Identification and Error Modeling of Automotive Speed Sensors
Figure 7: Error variance analysis on passive wheel speed sensor data gathered throughout three different trajectories.
Fig. 7 shows the results for the data collected in the cities of Kingston and Toronto that were used to fit the error variance. Upon initial data inspection, the quantity of interest that was found to impact measurement accuracy was the speed of travel. This is a clear indication that the sensor used is $passive$ in its nature. Moreover, the reason why the 0-4 $km/h$ speed interval was omitted in the variance analysis is simply due to the fact that PWSS may not report speeds lower than around 3 $km/h$ as shown in Fig. 8.
@@
The vehicular forward speed measurements were directly compared against the resultant velocity computed from the high-end tactical-grade integrated navigation solution. After further data analysis, a 2nd degree exponential function was selected in the data fitting process as it provides a flexible and efficient way to model the data accounting for non-linearities. The 2nd degree exponential function takes the form of:
@@
where the coefficients $a$, $b$, $c$, and $d$, are determined using an optimization algorithm. In this case, the obtained function is $\sigma^{2}_{PWSS}(v_{f})=0.3095e^{-0.1241v_{f}}+0.1477e^{-0.0009149v_{f}}$ and the coefficients are determined with 95% confidence bounds. This means that there is a 95% probability that the actual values of the coefficients fall within the bounds of the confidence intervals which corresponds to ±$\pm 3\sigma$. The obtained function provides a good representation of the data and can be used for further analysis or predictions.
Figure 9: PDF in terms of speed intervals for three real road trajectories, two of which are in Toronto, Ontario, Canada.
Driving patterns vary significantly depending on the environment, with vehicles typically traveling slower in dense urban areas and downtown cores compared to highways and other parts of the road network. Fig. 9 illustrates this trend, showing that the probability of a car traveling at speeds between 0–4 $km/h$ is 35–45% for trajectories T19.1 and K19.1, which traverse urban and downtown cores. In contrast, for T19.2, which includes a significant portion of highways, this probability drops below 30%. In dense, GNSS-denied urban areas, evaluating perception systems for speed estimation becomes particularly relevant.
@@ IV-B2 OBD-II-derived speed vs Stereo VO-derived speed
On the other hand, Fig. 12 illustrates the challenges faced in low-traffic environments dominated by static objects, which significantly affect the approach’s effectiveness. In such scenarios, Stereo VO generates noisier and less accurate speed estimates compared to OBD-II-derived speeds. The performance of Stereo VO-SS-OR deteriorates further at higher travel speeds, primarily due to baseline discrepancies
@@
between frames combined with framerate limitations. Additionally, the abundance of parked cars (as highlighted in the figure), often being misclassified as dynamic within the SS-OR routine, disrupts temporal feature tracking, and impairs accurate ego-motion estimation. Given these differences, implementing a fusion scheme is essential to selectively weight measurements based on WSS speed data, enabling a more accurate estimation of forward speed. Additionally, this highlights the critical need for integrating object tracking or supplementary sensors, such as RADAR sensors to improve stationary and dynamic object classification, enhancing the overall performance of the SS-OR framework.
@@ V Conclusion
In conclusion, this paper sheds light on a critical yet underexplored aspect of navigation and instrumentation: the diverse sensors and methods of deriving automotive speed measurements from the OEM perspective. By identifying and analyzing these approaches, we demonstrated the strengths and limitations of using OBD-II-derived speed data for land vehicle navigation. Through experiments conducted on three extended trajectories, we identified the wheel speed sensor type employed and showcased the advantages of using alternative ADAS sensors, such as front-facing stereo cameras, for speed estimation.
While each sensor modality has its own strengths and limitations, our findings emphasize the necessity of a fusion engine to combine data sources effectively, ensuring a robust integration with INS or RISS systems. To facilitate universal plug-and-play functionality, it would be beneficial to gather OBD-II-derived speed data from a wide range of vehicles, including those equipped with active wheel speed sensors. This approach would enable the development of more accurate error profiles and an adaptive system that enhances the fusion process with low-cost exteroceptive sensors.
Looking ahead, for more sophisticated and car-specific systems, reverse engineering the CAN bus to access individual wheel speed data directly presents a promising avenue. This approach has the potential to significantly enhance the accuracy of speed estimation, ultimately leading to substantial improvements in overall navigation system performance.

### v3

@@ II Related Work
In addition to built-in vehicular speed sensing and DMIs, alternative systems have been explored for specialized ground vehicle testing. Magnetic Rotary Encoders (MREs) [7] and Non-contact Optical Sensors [8, 9] are examples used in extreme conditions like snow, ice, or wet surfaces, or in harsh environments with dust, moisture, extreme temperatures, and mechanical shocks. Despite their robustness and precision, these niche applications and high costs limit their practical use in everyday automotive systems.
@@
[b] Aspect OBD-II Standard CAN Bus Purpose Diagnostics and emissions monitoring Real-time communication between ECUs Standardization Standardized PIDs (e.g., 0x0D) Manufacturer-specific (i.e., CAN ID varies) Data Access Limited to subset of vehicle parameters Full access to all CAN messages Communication Model Request-Response Broadcast (continuous streaming) Speed of Access Slower, query-based Faster, real-time Physical Access OBD-II port OBD-II port or direct CAN wiring
@@
  • PID: Parameter ID as defined in SAE J1979.
  • CAN remains the most adopted in-vehicle network; Ethernet (802.3bw) offers similar features and higher data rates for emerging applications.
In addition to the often omitted derivation process, the navigation and instrumentation literature offers limited clarity on the specific sensors automotive suppliers use for speed estimation. Researchers often overlook the type of speed sensors employed in their test vehicles, commonly employ the general term odometer which covers a wide variety of instruments, and frequently rely on traditional Inertial Measurement Unit (IMU) and GNSS error modeling techniques for sensor fusion [4, 12, 13, 14]. While these approaches have been widely used, they may not guarantee optimal estimation of vehicle speed. Therefore, identifying and characterizing vehicular speed sensors is critical for designing robust and effective navigation systems.
Fig. 2 categorizes vehicle speed-sensing technologies into two primary groups: built-in diagnostic data and alternative sensors and systems. The first group includes OBD-II-derived speed measurements and CAN bus-derived data from Vehicle Speed Sensors (VSS) and WSS, both of which are widely integrated into vehicles’ onboard systems. The second group consists of external sensors, such as non-contact optical sensors, DMIs, and perception-based systems like cameras, LiDARs, and RADARs. Cameras and RADARs, in particular, are increasingly utilized not only for ADAS but also for forward velocity estimation. Recent studies have demonstrated the effectiveness of cameras for ego-motion tracking [15] and RADARs for ego-velocity estimation [16]. These perception-based approaches offer practical and cost-effective alternatives, particularly in scenarios where traditional automotive speed sensors are limited or unavailable.
@@ III OBD-II Vehicle Speed Derivation
Modeling errors in consumer-level OBD-II-derived speed measurements presents a significant challenge due to the restricted data available through the vehicle’s regular OBD-II protocol. Generally, the OBD-II standard provides data on the overall vehicle speed rather than the speed of each individual wheel. Additionally, the speed data is often in integer format, which lacks the resolution of decimal points found in double or float types commonly available in CAN bus systems, leading to less precise measurements. When querying the OBD-II system for vehicle speed, a command is sent, and the system responds with encoded data. Below is an example illustrating this process:
@@
The command 010D represents a request for the current vehicle speed (PID 0x0D). The response 41 0D 3C indicates the vehicle speed as 0x3C, which is 60 $km/h$ when converted to decimal [^1] Automakers may configure ECUs to report diagnostic data in region-specific units. For example, vehicle speed may be in miles per hour ($mph$) for the USA or in kilometers per hour ($km/h$) for Canada and Europe.. In other words, the reported speed through standard OBD-II interfacing is:
@@ III-A Transmission-based Vehicle Speed Determination
The estimation of vehicle speed ($V_{\text{v}}$) from transmission data relies on the relationship between engine speed in RPM and drivetrain parameters, especially gear ratio. The gear ratio governs the conversion of engine speed (input) to wheel speed (output). For example, in Fig. 3, a driving gear with 8 teeth and a driven gear with 16 teeth creates a 2:1 ratio, where the output rotates at half the input speed. This adjustment is crucial for controlling speed and torque. The gearbox alters engine speed based on the engaged gear ratio ($i_{x}$), which changes with each gear shift, while the differential uses its gear ratio ($i_{0}$) to balance wheel speeds, critical for smooth turning.
@@
$V_{\text{v}}^{\text{OBD}}=\left\lfloor\left({\frac{N_{e}\cdot\pi\cdot r_{w}}{30\cdot i_{x}\cdot i_{0}}}\right)\right\rceil,$ (2)
$N_{e}$ represents the engine speed in RPM, typically measured by a crankshaft position (CKP) sensor or an engine speed sensor. The wheel radius ($r_{w}$), measured in meters ($m$), converts the rotational speed of the wheels into linear speed. This calculation assumes a slip-free mechanical connection between the engine and wheels, ensuring efficient power transfer and accurate speed estimation.
@@ III-B ABS-based Vehicle Speed Determination
where $V_{\text{w}}$ denotes the linear speed of an individual wheel ($m/s$), $N_{p}$ is the number of pulses detected during the sampling interval $\Delta t$ ($s$), and $N_{t}$ is the total number of teeth on the reluctor ring. To determine the vehicle’s overall speed, the ECU averages the rotational speeds of all the individual wheels. This average is then used for various control and monitoring functions within the vehicle’s ABS and electronic stability control (ESC) systems.
Compared to the transmission-based system, ABS-based speed estimation offers greater robustness and accuracy while involving fewer parameters. Thus, mathematically, the actual ABS wheel speed reported through the regular OBD-II protocol is computed as follows:
$V_{\text{v}}^{\text{OBD}}=\left\lfloor\left(\frac{V_{\text{w}_{1}}+V_{\text{w}_{2}}+V_{\text{w}_{3}}+V_{\text{w}_{4}}}{4}\right)\right\rceil,$ (4)
where $V_{\text{w}_{i}}$ denotes the velocity of the $i^{th}$ wheel ($m/s$), assuming each wheel is equipped with a WSS. To avoid reliance on automotive OEM collaboration or specialized tools (e.g., CAN bus reverse engineering) for accessing detailed sensor information in test vehicles, navigation and mapping entities often rely on carefully installed and calibrated DMIs.
@@ III-C Wheel Speed Sensor Types
When dealing with WSS, limited information can be found in the navigation and instrumentation literature related to the type of WSS automotive suppliers use and their corresponding accuracy. Within the automotive industry, two primary types of WSS are currently utilized, each possessing unique advantages and disadvantages. These are referred to as $passive$ and $active$ wheel speed sensors, respectively. The operational mechanisms of these sensors exceed the scope of this paper, however, their main differences are shown in Table II.
@@ IV-A Experimental Setup
The experiments were conducted using an ABS-equipped Toyota Sienna minivan, with the data logging setup outlined in Fig. 6. A detailed dataset description is provided in [18], and supporting materials are publicly available [^2] GitHub repository: https://github.com/hanymragab/plans2025-resources.
@@ IV-B1 Identification and Error Modeling of Automotive Speed Sensors
To identify the type of WSS, we first examined the speed profiles for values below 3 $km/h$. If such readings were present, the sensor was classified as active; otherwise, it was considered passive. This approach aligns with the known behavior of passive WSS, which typically do not report speeds below this threshold. A sample speed profile from one of the test trajectories (Fig. 7) shows no data in the low-speed range, indicating the use of passive sensors. Given that passive WSS accuracy improves with increasing speed, the error variance analysis was conducted as a function of travel speed. Fig. 8 presents results from Kingston and Toronto, which reflect this trend and were used to fit the variance model.
Figure 8: Error variance analysis of passive wheel speed sensor data from three trajectories: ‘T’ and ‘K’ indicate Toronto and Kingston, respectively, while ‘19’ refers to data collected in 2019.
The vehicular speed measurements were directly compared against the resultant velocity computed from the high-end tactical-grade integrated navigation solution. After further data analysis, a 2nd degree exponential function was selected in the data fitting process as it provides a flexible and efficient way to model the data accounting for non-linearities. The 2nd degree exponential function takes the form of:
@@
where the coefficients $a$, $b$, $c$, and $d$, are determined using an optimization algorithm. In this case, the obtained function is $\sigma^{2}_{PWSS}(v_{f})=0.3095e^{-0.1241v_{f}}+0.1477e^{-0.0009149v_{f}}$ and the coefficients are determined with 95% confidence bounds. This means that there is a 95% probability that the actual values of the coefficients fall within the bounds of the confidence intervals which corresponds to ±$\pm 2\sigma$. The obtained function provides a good representation of the data and can be used for further analysis or predictions.
Figure 9: Histogram-based Probability Mass Function (PMF) in terms of speed intervals for three real road trajectories, two of which are in Toronto.
Driving patterns vary by environment, with vehicles typically moving slower in dense urban areas and downtown cores than on highways. As shown in Fig. 9, the histogram-based probability mass function (PMF) shows that speeds between 0–4 km/h account for 35–45% of samples in urban trajectories T19.1 and K19.1. In contrast, for T19.2, which includes highway segments, this falls below 30%. These trends highlight the need to assess perception-based speed estimation in GNSS-challenging urban settings, where passive WSS may fail to capture low-speed motion.
@@ IV-B2 OBD-II-derived speed vs Stereo VO-derived speed
On the other hand, Fig. 12 illustrates the challenges faced in low-traffic environments dominated by static objects, which significantly affect the approach’s effectiveness. In such scenarios, Stereo VO generates noisier and less accurate speed estimates compared to OBD-II-derived speeds. The performance of Stereo VO-SS-OR deteriorates further at higher travel speeds, primarily due to baseline discrepancies between frames combined with framerate limitations.
Additionally, the abundance of parked cars (as highlighted in the figure), often being misclassified as dynamic within the SS-OR routine, disrupts temporal feature tracking, and impairs accurate ego-motion estimation. Given these differences, implementing a fusion scheme is essential to selectively weight measurements based on WSS speed data, enabling a more accurate estimation of forward speed. Additionally, this highlights the critical need for integrating object tracking or supplementary sensors, such as RADAR sensors to improve stationary and dynamic object classification, enhancing the overall performance of the SS-OR framework.
@@ V Conclusion
In conclusion, this paper sheds light on a critical yet underexplored aspect of navigation and instrumentation: the diverse sensors and methods for deriving automotive speed measurements from the OEM perspective. By identifying and analyzing these approaches, we demonstrated the strengths and limitations of using OBD-II-derived speed data for land vehicle navigation. Through experiments conducted on three long real road trajectories, we identified the WSS type employed and showcased the advantages of using alternative ADAS sensors, such as front-facing stereo cameras, for speed estimation.
While each sensor modality offers unique advantages and limitations, our findings underscore the importance of a fusion engine capable of reliably integrating diverse data sources to ensure robust and resilient navigation performance. Achieving universal plug-and-play capability would require collecting OBD-II-derived speed data across a broad spectrum of vehicles, including those equipped with active wheel speed sensors. This expanded dataset would support the development of more accurate and vehicle-specific error profiles, enabling a more adaptive and generalizable fusion framework that leverages low-cost exteroceptive sensors.
Looking ahead, for more advanced and vehicle-specific systems, reverse engineering in-vehicle communication protocols to access individual wheel speed data directly presents a promising direction. This approach can significantly improve the accuracy of speed estimation and, in turn, enhance the overall performance of the integrated navigation system.